# Fading in the Flow: Suppression of cold gas growth in expanding galactic outflows

### v2

@@ Appendix B Cloud surface & scaling relations
Figure 16: Evolution of the surface area (in units of the initial cloud surface area) of the radiatively cooling turbulent boundary layer, identified as the bounding surface of all gas below a temperature of $2T_{\rm cl}=8\times 10^{4}\ \rm K$, as a function of cloud volume (upper panel) the distance (lower panel) travelled by the center of mass of the cloud (normalized by the initial cloud distance from the center of the CC85 wind). Along each of the curves, the diamonds and stars have the same meaning as in Fig. 9. The surface area of the clouds, after entrainment, follows $A_{\rm cl}\propto V_{\rm cl}^{2/3}$ as indicated by the yellow band in the upper panel. As discussed in Section 4.1.1, one gets $A_{\rm cl}\propto d_{\rm cl}^{3\gamma-2}\sim d_{\rm cl}^{3};\gamma=5/3$, which is approximately followed by the largest clouds (yellow band in the lower panel) that do not grow and experience as much orthogonal expansion compared to their initial size as the smaller clouds (cf. Fig. 9). While for small clouds, relative expansion is important and $m_{\rm cl}\propto d_{\rm cl}^{23/12}\approx d_{\rm cl}^{2}$ (cf. Fig. 10), which implies $A_{\rm cl}\propto d_{\rm cl}^{3(\gamma-1/2)}\sim d_{\rm cl}^{3.5};\gamma=5/3$ as indicated by the gray band in the lower panel, albeit the difference in the slopes between the yellow and the gray bands is small (compare with Fig. 10 where the difference in the slopes is large).
We investigate how the surface area scales with the volume of the cold gas in a CC85 wind. From geometric considerations, the cloud surface area $A_{\rm cl}$ should follow $A_{\rm cl}\propto V_{\rm cl}^{2/3}$, where $V_{\rm cl}$ is the volume of the cold cloud. This scaling relation is highlighted in a yellow band in the upper panel of Fig. 16, which shows that as clouds start to grow with strong radiative cooling, their surface area follows the simple geometric (non-fractal) scaling relation.
Figure 17: Evolution of the inflow velocity (in units of the wind speed at the initial cloud position) across the surface of the radiatively cooling turbulent boundary layer, identified in the same way as in Fig. 16, as a function of the distance travelled by its center of mass (normalized by its initial distance from the center of the CC85 wind). The inflow velocity $v_{\rm in}$ has been calculated as the average mass-weighted velocity along the local normals used to tesselate the iso-surface of temperature at $8\times 10^{4}\ \rm K$ bounding the cloud. Albeit with a lot of scatter, $v_{\rm in}$ scales with the position of the center-of-mass $d$ of the clouds as $v_{\rm in}\propto d_{\rm cl}^{-\gamma/2}\sim d_{\rm cl}^{-5/6};\gamma=5/3$, as indicated by the yellow band.
@@
We also calculate the (mass) average inflow velocity $v_{\rm in}$ across the surface of the cold gas at different instants as a function of the center-of-mass of the cloud $d_{\rm cl}$ with respect to the initial position. To calculate this inflow velocity, we identify the iso-suface bounding all gas below $2T_{\rm cl}=8\times 10^{4}\ \rm K$ and calculate the local normal for every tile in the tesselation of the iso-surface. The inflow velocity $v_{\rm in}$ is the mass averaged velocity along the local normal for all the tiles bounding the iso-surface (see Abruzzo et al. 2024 for details). This inflow velocity $v_{\rm in}$, although employing different definitions, was previously shown by Fielding et al. 2020; Gronke & Oh 2020; Tan et al. 2021 to follow the scaling relation $v_{\rm in}\propto t_{\rm cool,cl}^{-1/4}$ in the strong cooling regime, where $t_{\rm cool,cl}$ is the cooling time of the cold gas – the fastest cooling phase. Using Eq. 3.1 and the arguments presented in Section 3.3 (that clouds are at $T_{\rm cl}$ and locally isobaric), we get $t_{\rm cool,cl}\propto d_{\rm cl}^{2\gamma}$. Therefore, the inflow velocity $v_{\rm in}$ follows $v_{\rm in}\propto d_{\rm cl}^{-\gamma/2}$, as indicated by the yellow band in Fig. 17. There is a significant scatter in the simulation data and this trend is only crudely followed by the clouds. The dependence of $R_{\rm cl,\perp}$ ($\sim d$ (see Section 4.1.1) introduces a weak additional dependence of $v_{\rm in}$ ($\sim R_{\rm cl,\perp}^{1/4}t_{\rm cool,cl}^{-1/4}\sim d^{1/4-\gamma/2}\approx d^{-\gamma/2};\gamma=5/3$) on the distance travelled by the cloud, which is even shallower but difficult to tell apart due to the noisy nature of $v_{\rm in}$.
Appendix C Clouds in an anisotropically expanding box
@@
Gronke & Oh 2020 (GO20 hereafter) employed a model of a homogeneous anisotropically expanding computational domain to approximate the evolution of a cloud in a CC85 wind. The cloud is assumed to be much smaller than its from the wind center. In addition, the entire cloud is assumed to experience a uniform wind condition, which only changes over time as the cloud moves down the wind. We re-simulated the cloud-crushing problem with the anisotropic expansion model used by GO20 in PLUTO for our fiducial parameters and ran it up to $30\ t_{\rm cc,ini}$. The extent of the simulation domain is $250R_{\rm cl}$ in the direction of the wind and $22R_{\rm cl}$ in directions orthogonal to it. Along with that, we utilize the frame boost method from Kanjilal et al. 2021 to track the cloud.
@@ Appendix D Convergence
Figure 19: Slice plots of the temeprature field comparing cloud-crushing simulations in a CC85 wind at four different resolutions – $R_{\rm cl}/d_{\rm cell}|_{\rm ini}=(8,16,32,64)$ at time $t/t_{\rm cc,ini}=6$. As expected, on increasing resolution, mixing at smaller scales is better resolved. For these simulations, a domain with an extent of about $250\ R_{\rm cl}$ in the radial direction and $22\ R_{\rm cl}$ at the initial position of the cloud in the othogonal direction had been considered, with the exception of the simulation with the highest resolution of $R_{\rm cl}/d_{\rm cell}|_{\rm ini}=64$, where the radial extent of the box has been reduced to $100\ R_{\rm cl}$ due to computational limitations. Follow the link here to see a video of the full evolution.
@@
Figure 20: Comparison of the evolution of cold cloud mass (left panel) and distance travelled by the center of mass of the cold gas (right panel) in CC85 wind from our simulation at different resolutions using the fiducial parameter $t_{\rm cool,mix}/t_{\rm cc}|_{\rm ini}=0.2$ (cf. Table 1). The resolution marked in the legend denotes the number of cells by which the cloud radius is resolved at the initial location of the cloud; a higher value corresponds to a higher resolution. The highest resolution curve with $R_{\rm cl}/d_{\rm cell}|_{\rm ini}=64$ has not been considered beyond $6\ t{\rm cc,ini}$ as a significant amount of cold gas leaves the computational domain due to a smaller domain size compared to the simulations at lower resolutions. The highest resolution curve $R_{\rm cl}/d_{\rm cell}|_{\rm ini}=64$ is confined within the $R_{\rm cl}/d_{\rm cell}|_{\rm ini}=16$ and $R_{\rm cl}/d_{\rm cell}|_{\rm ini}=32$ curves, indicating that the cold mass growth is converged. The mass estimated at different resolutions agree with one another within a factor of a few, and their difference diminishes as the simulation resolution increases (compare, for example, the difference between the $R_{\rm cl}/d_{\rm cell}|_{\rm ini}=32$ and $R_{\rm cl}/d_{\rm cell}|_{\rm ini}=64$) indicating convergence in cloud mass. Increased mixing at high resolution causes increased momentum transfer from the hot to the cold phase causing the cloud to move faster in the wind at higher resolution.
@@
We demonstrate results from our convergence test on the evolution of the cold cloud mass in our simulation for the fiducial parameter, $t_{\rm cool,mix}/t_{\rm cc}|_{\rm ini}=0.2$ (cf. Table 1). For all simulations in our convergence test, a domain with an extent of about $250\ R_{\rm cl}$ in the radial direction and $22\ R_{\rm cl}$ at the initial position of the cloud in the orthogonal directions had been considered. The only exception is the simulation with the highest resolution of $R_{\rm cl}/d_{\rm cell}|_{\rm ini}=64$, where the radial extent is reduced to $100\ R_{\rm cl}$ due to computational limitations. We find that with higher resolution (increasing the value of $R_{\rm cl}/d_{\rm cell}|_{\rm ini}$ at the initial position of the cloud), the cold mass in the cloud is higher at all times. However, the results across different resolutions agree within a factor of a few in the time window of our analysis. At high resolutions, the evolution of the cold mass tends to converge, although the clouds differ considerably in their morphology. The qualitative picture of cloud growth due to strong radiative cooling is well captured across all resolutions. The increased amount of cold mass with increasing resolution is not surprising, as we are able to capture the formation and evolution of smaller cold clumps when the resolution is higher. Dissipation in our simulations operates at the grid resolution, and there are no explicit dissipative terms. Consequently, with increased resolution, the grid is able to capture the mixing and cooling till smaller scales. This increased resolution causes an increase in mixing and cooling, resulting in a slightly higher transfer of mass and momentum from the hot to the cold phase, as seen in the position of the center of mass of the cloud in the right panel of Fig. 20. Therefore, the cloud travels faster and gets quickly entrained at higher resolution. A similar trend has also been previously found in our earlier works on cloud-crushing (cf. Fig. B1 in Kanjilal et al. 2021). However, CC85 wind brings an additional complication by the faster travel of the cloud material – the wind condition depends strongly on the instantaneous position of the cloud (for example, $T_{\rm wind}\sim d^{-2(\gamma-1)}$), which can lead to significant differences in cloud properties across simulations at different resolutions as time progresses.
Appendix E Single-cloud model from FB22
$\rm\frac{d}{dt}\begin{bmatrix}d_{\rm cl}\\ v_{\rm cl}\\ M_{\rm cl}\end{bmatrix}=\begin{bmatrix}v_{\rm cl}\\ v_{\rm rel}\left(\dot{M}_{\rm cl,grow}/M_{\rm cl}\right)+(3/8)C_{\rm drag}v_{\rm rel}^{2}/\left(\chi R_{\rm cl}\right)\\ \dot{M}_{\rm cl,0}\left(\xi^{\alpha}-f_{\rm mix}/f_{\rm cool}\right)\end{bmatrix}.$ (E.1)
[…]

### v3

@@ Appendix B Cloud surface & scaling relations
Figure 17: Evolution of the surface area (in units of the initial cloud surface area) of the radiatively cooling turbulent boundary layer, identified as the bounding surface of all gas below a temperature of $2T_{\rm cl}=8\times 10^{4}\ \rm K$, as a function of cloud volume (upper panel) the distance (lower panel) travelled by the centre of mass of the cloud (normalised by the initial cloud distance from the centre of the CC85 wind). Along each of the curves, the diamonds and stars have the same meaning as in Fig. 10. The surface area of the clouds, after entrainment (beyond the star marks), follows $A_{\rm cl}\propto V_{\rm cl}^{2/3}$ as indicated by the yellow band in the upper panel. However, prior to entrainment, the area of the boundary layer (formed by shear induced instabilities) increases much steeply with volume, as indicated by the cyan band in the upper panel which marks $A_{\rm cl}\propto V_{\rm cl}^{5/6}$, due to radiative cooling which gives a fractal nature to the cloud surface (Fielding et al. 2020). Although the differences in the two slopes in the area-volume relation is small, there is a slight flattening of the slope beyond the star marks in the upper panel. As discussed in Section 4.1.1, one gets $A_{\rm cl}\propto d_{\rm cl}^{3\gamma-2}\sim d_{\rm cl}^{3};\gamma=5/3$, which is approximately followed by the largest clouds (yellow band in the lower panel) that do not experience as much orthogonal expansion compared to their initial size as the smaller clouds (cf. Fig. 10). While for small clouds, this relative expansion is important and $m_{\rm cl}\propto d_{\rm cl}^{23/12}\approx d_{\rm cl}^{2}$ (cf. Fig. 11), which implies $A_{\rm cl}\propto d_{\rm cl}^{3(\gamma-1/2)}\sim d_{\rm cl}^{3.5};\gamma=5/3$ as indicated by the cyan band in the lower panel, albeit the difference in the slopes between the yellow and the cyan bands is small (compare with Fig. 11 where the difference in the slopes is large).
Figure 18: Evolution of the extent of a cloud $R_{\rm cl,\parallel}$ in the direction of the wind (in units of the initial cloud radius) as a function of the distance travelled by the centre of mass of the cloud (normalised by the initial cloud distance from the centre of the CC85 wind). The length of the cloud has been calculated as the difference between the $60$ and 55 percentile values of the position of the cloud material along the axis passing through the centre of the cloud parallel to the wind. Along each of the curves, the diamonds and stars have the same meaning as in Fig. 10. For clouds that are entrained (beyond the stars), the length of the cloud in the wind direction approximately saturates to a constant. On the other hand, the largest clouds which do not get entrained in their journey down the wind keep growing in their extent in the wind direction proportional to the distance they travel as indicated by the yellow band. This evolution can be compared with the orthogonal extent of the cloud (cf. Fig. 10) which shows a different scaling for the two types of clouds categorized according to their size.
We investigate how the surface area scales with the volume of the cold gas in a CC85 wind. To evaluate the cloud surface area, we first identify the all the cells in the simulation domain with temperature less than $8\times 10^{4}\ \rm K$ within a region where the CC85 wind temperature is greater than $9\times 10^{4}\ \rm K$ using the VTK ‘Threshold’ filter that comes as a part of the Paraview data analysis/visualisation software. We then isolate the outer surface of all these ‘cold’ cells by selecting only the cell faces that belong to one cell and are not shared by any other cells using the ‘Extract Surface’ VTK filter also available in Paraview. Additionally to calculate the inflow velocity, we interpolate the fields to the face centres of the faces tessellating this outer surface and calculate the surface area of every face as well as their respective surface normals. [^10] Sanity checks on the initial spherical cloud using these VTK filters give consistent results for the cloud volume and surface area.
Once the cloud is sufficiently entrained, from geometric considerations, we propose that the cloud surface area $A_{\rm cl}$ to follow $A_{\rm cl}\propto V_{\rm cl}^{2/3}$, where $V_{\rm cl}$ is the volume of the cold cloud. This scaling relation is highlighted in a yellow band in the upper panel of Fig. 17, which shows that as clouds start to grow with strong radiative cooling, their surface area follows the simple geometric (non-fractal) scaling relation. However, initially, the cloud area increases much rapidly with volume. For the largest clouds that do not get sufficiently entrained the surface area is dominated by shear-driven turbulent mixing which radiatively cools and has a fractal nature (Fielding et al. 2020), where $A_{\rm cl}\propto V_{\rm cl}^{5/6}$ (fractal dimension of $2.5$) – a steeper power law dependence between surface area and cloud volume. Although the differences in the slope are small and there is a lot of scatter in the data, a gradual flattening of the scaling relation post entrainment between surface area and cloud volume is evident. Large clouds that do not get entrained also behave differently in terms of their size along the wind which keeps growing with the distance they travel compared to the smaller clouds, whose extent along the wind saturates (cf. Fig. 18). Further investigation is needed to determine the influence of simulation resolution on the fractal surface geometry of clouds and its potential consequences for the emergent scaling relation.
@@
We also calculate the (mass) average inflow velocity $v_{\rm in}$ across the surface of the cold gas at different instants as a function of the centre-of-mass of the cloud $d_{\rm cl}$ with respect to the initial position. To calculate this inflow velocity, we identify the iso-surface bounding all gas below $2T_{\rm cl}=8\times 10^{4}\ \rm K$ and calculate the local normal for every tile in the tessellation of the iso-surface. The inflow velocity $v_{\rm in}$ is the mass averaged velocity along the local normal for all the tiles bounding the iso-surface (see Abruzzo et al. 2024 for details). This inflow velocity $v_{\rm in}$, although employing different definitions, was previously shown by Fielding et al. 2020; Gronke & Oh 2020; Tan et al. 2021 to follow the scaling relation $v_{\rm in}\propto t_{\rm cool,cl}^{-1/4}$ in the strong cooling regime, where $t_{\rm cool,cl}$ is the cooling time of the cold gas – the fastest cooling phase. Using Eq. 3.1 and the arguments presented in Section 3.2 (that clouds are at $T_{\rm cl}$ and locally isobaric), we get $t_{\rm cool,cl}\propto d_{\rm cl}^{2\gamma}$. Therefore, the inflow velocity $v_{\rm in}$ follows $v_{\rm in}\propto d_{\rm cl}^{-\gamma/2}$, as indicated by the yellow band in Fig. 19. There is a significant scatter in the simulation data and this trend is only crudely followed by the clouds. The dependence of $R_{\rm cl,\perp}$ ($\sim d$ (see Section 4.1.1) introduces a weak additional dependence of $v_{\rm in}$ ($\sim R_{\rm cl,\perp}^{1/4}t_{\rm cool,cl}^{-1/4}\sim d^{1/4-\gamma/2}\approx d^{-\gamma/2};\gamma=5/3$) on the distance travelled by the cloud, which is even shallower but difficult to tell apart due to the noisy nature of $v_{\rm in}$.
Figure 19: Evolution of the inflow velocity (in units of the wind speed at the initial cloud position) across the surface of the radiatively cooling turbulent boundary layer, identified in the same way as in Fig. 17, as a function of the distance travelled by its centre of mass (normalised by its initial distance from the centre of the CC85 wind). The inflow velocity $v_{\rm in}$ has been calculated as the average mass-weighted velocity along the local normals used to tessellate the iso-surface of temperature at $8\times 10^{4}\ \rm K$ bounding the cloud. Albeit with a lot of scatter, $v_{\rm in}$ scales with the position of the centre-of-mass $d$ of the clouds as $v_{\rm in}\propto d_{\rm cl}^{-\gamma/2}\sim d_{\rm cl}^{-5/6};\gamma=5/3$, as indicated by the yellow band.
Appendix C Single-cloud model from FB22
Figure 20: Comparison of the distance travelled by the centre of mass of the cloud from our cloud-crushing simulation with that predicted from the FB22 single-cloud model for ‘vanilla’ cloud-crushing (in uniform wind) and for cloud-crushing in a CC85 wind (see Section 4.1.2 for details). The dashed lines are the simulation results for ‘vanilla’ cloud-crushing (in yellow) and cloud-crushing in a CC85 background (in teal). As a specific example which would clearly demonstrate departure from model prediction, we choose the simulation result with $t_{\rm cool,mix}/t_{\rm cc}|_{\rm ini}=0.8$ and $\mathcal{M}=1.496$ (cf. Table 1; not the fiducial parameter set). The corresponding model predictions are shown with solid lines in the same colour as their dashed counterparts. Although Fig. 13 shows that the mass of clouds has a large variation between the original parameter set from FB22 and our set of parameters (described in Section 4.1.2), we find that the distance the cloud travels with time differs from our parameters (solid dark lines) and the original one (thick, solid light-coloured curves) by less than $1\%$.
$\rm\frac{d}{dt}\begin{bmatrix}d_{\rm cl}\\ v_{\rm cl}\\ M_{\rm cl}\end{bmatrix}=\begin{bmatrix}v_{\rm cl}\\ v_{\rm rel}\left(\frac{\dot{M}_{\rm cl,grow}}{M_{\rm cl}}\right)+\frac{3}{8}C_{\rm drag}\frac{v_{\rm rel}^{2}}{\chi R_{\rm cl}}\\ \dot{M}_{\rm cl,0}\left(\frac{v_{\rm in}}{f_{\rm turb}|v_{\rm rel}|}-\frac{f_{\rm mix}}{f_{\rm cool}}\right)\end{bmatrix}$, (C.1)
where $v_{\rm in}$ is the inflow velocity across the turbulent boundary layer, where the ansatz $v_{\rm in}=f_{\rm turb}|v_{\rm rel}|\xi^{\alpha}$ is used in the original FB22 single-cloud model. We modify $v_{\rm in}$ and propose a ‘modified’ single-cloud model as represented in Eq. 4.1. This sets a floor value to $v_{\rm in}$ such that the model prediction can better match the trend in mass evolution of cold clouds from the simulation data (cf. Section 4.1.2).
[…]
In Fig. 20, we show a comparison of the model prediction of the distance travelled by the cloud using FB22 single-cloud model (with the original parameters as well as our modified one) with simulation results (similar to Fig. 13). Interestingly, we find that unlike evolution of cloud mass, the distance travelled by the cloud predicted by the model is almost unaffected (variation $\lesssim 1\%$) on changing from the original FB22 parameters to our. Both models predict that a cloud would travel a greater distance downstream in an expanding CC85 wind compared to a uniform wind – consistent with the trend found in the simulations. It is worth noting that we fit our model prescription to only match the mass evolution, while the distance evolution comes as a natural consequence of the chosen model. In addition, cloud mass $m_{\rm cl}\propto d_{\rm cl}^{3}$ in a uniform wind compared to $m_{\rm cl}\propto d_{\rm cl}^{2}$ (or $m_{\rm cl}\propto d_{\rm cl}$; cf. Fig. 11) in a CC85 wind causing clouds in uniform wind to grow more in mass even though they travel lesser distance downstream compared to clouds in an expanding wind.
Appendix D Clouds in an anisotropically expanding box
@@
Gronke & Oh 2020 (GO20 hereafter) employed a model of a homogeneous anisotropically expanding computational domain to approximate the evolution of a cloud in a CC85 wind. The cloud is assumed to be much smaller than its from the wind centre. In addition, the entire cloud is assumed to experience a uniform wind condition, which only changes over time as the cloud moves down the wind. We re-simulated the cloud-crushing problem with the anisotropic expansion model used by GO20 in PLUTO for our fiducial parameters and ran it up to $30\ t_{\rm cc,ini}$. The extent of the simulation domain is $250R_{\rm cl}$ in the direction of the wind and $22R_{\rm cl}$ in directions orthogonal to it. Along with that, we utilise the frame boost method from Kanjilal et al. 2021 to track the cloud.
@@ Appendix E Convergence
Figure 22: Slice plots of the temperature field comparing cloud-crushing simulations in a CC85 wind at four different resolutions – $R_{\rm cl}/d_{\rm cell}|_{\rm ini}=(8,16,32,64)$ at time $t/t_{\rm cc,ini}=6$. As expected, on increasing resolution, mixing at smaller scales is better resolved. For these simulations, a domain with an extent of about $250\ R_{\rm cl}$ in the radial direction and $22\ R_{\rm cl}$ at the initial position of the cloud in the orthogonal direction had been considered, with the exception of the simulation with the highest resolution of $R_{\rm cl}/d_{\rm cell}|_{\rm ini}=64$, where the radial extent of the box has been reduced to $100\ R_{\rm cl}$ due to computational limitations. Follow the link here to see a video of the full evolution.
@@
Figure 23: Comparison of the evolution of cold cloud mass $m_{\rm cl}(t)$ (upper left panel), its growth rate $\dot{m}_{\rm cl}(t)$ (lower left panel) and distance travelled by the centre of mass of the cold gas $d_{\rm cl}(t)$ (right panel) in CC85 wind from our simulation at different resolutions using the fiducial parameter $t_{\rm cool,mix}/t_{\rm cc}|_{\rm ini}=0.2$ (cf. Table 1). The resolution marked in the legend denotes the number of cells by which the cloud radius is resolved at the initial location of the cloud; a higher value corresponds to a higher resolution. The highest resolution curve with $R_{\rm cl}/d_{\rm cell}|_{\rm ini}=64$ has not been considered beyond $6\ t_{\rm cc,ini}$ as a significant amount of cold gas leaves the computational domain due to a smaller domain size compared to the simulations at lower resolutions. The highest resolution curve $R_{\rm cl}/d_{\rm cell}|_{\rm ini}=64$ is confined within the $R_{\rm cl}/d_{\rm cell}|_{\rm ini}=16$ and $R_{\rm cl}/d_{\rm cell}|_{\rm ini}=32$ curves, indicating that the cold mass growth is converged. The mass estimated at different resolutions agree with one another within a factor of a few, and their difference diminishes as the simulation resolution increases (compare, for example, the difference between the $R_{\rm cl}/d_{\rm cell}|_{\rm ini}=32$ and $R_{\rm cl}/d_{\rm cell}|_{\rm ini}=64$) indicating convergence in cloud mass. In fact, the mass growth rates $\dot{m}_{\rm cl}$ are even better converged across different resolutions. Increased mixing at high resolution causes increased momentum transfer from the hot to the cold phase causing the cloud to move faster in the wind at higher resolution.
Figure 24: Evolution of the cloud profile at a resolution $R_{\rm cl}/d_{\rm cell}|_{\rm ini}=16$; one level higher than our fiducial resolution with the same fiducial parameters (cf. Table 1) as in Fig. 7. The clouds maintain local pressure equilibrium with the background wind and falls steeply in their density contrast with respect to the background. In terms of these general features, the cloud properties remain unchanged across resolution demonstrating the robustness of our simulation results and the analytic scaling relations presented in Section 3. Although the cloud pressure profiles track the local wind pressure, it is always slightly under-pressurized with respect to the ambient medium which is sustained even after increasing resolution – tentative evidence of a local cooling flow around these clouds (Dutta et al. 2022), rather than an artifact of low resolution.
@@
We demonstrate results from our convergence test on the evolution of the cold cloud mass in our simulation for the fiducial parameter, $t_{\rm cool,mix}/t_{\rm cc}|_{\rm ini}=0.2$ (cf. Table 1). For all simulations in our convergence test, a domain with an extent of about $250\ R_{\rm cl}$ in the radial direction and $22\ R_{\rm cl}$ at the initial position of the cloud in the orthogonal directions had been considered. The only exception is the simulation with the highest resolution of $R_{\rm cl}/d_{\rm cell}|_{\rm ini}=64$, where the simulation domain is reduced to $100\ R_{\rm cl}$ in the radial direction due to computational limitations. We find that with higher resolution (increasing the value of $R_{\rm cl}/d_{\rm cell}|_{\rm ini}$ at the initial position of the cloud), the cold mass in the cloud is mostly higher – except for the $R_{\rm cl}/d_{\rm cell}|_{\rm ini}=32$ and $64$ simulations where $m_{\rm cl}(t)$ seems to agree. Also with the other simulations, the results across different resolutions agree within a factor of a few in $m_{\rm cl}(t)$ and $\dot{m}_{\rm cl}(t)$ (see left panels in Fig. 23) in the time window of our analysis. At high resolutions, the evolution of the cold mass tends to converge, although the clouds differ considerably in their morphology. The qualitative picture of cloud growth due to mixing and strong radiative cooling is well captured across all resolutions.
The convergence of cold gas mass growth is not surprising – and has been studied extensively in prior cloud-crushing (e.g. Gronke & Oh, 2020) and high resolution shearing layer simulations (e.g. Tan et al., 2021). Although we do not explicitly model thermal conduction, therefore, the thermal diffusion across a cold gas front is purely numerical in our simulations. Nevertheless, this is sub-dominant compared to turbulent diffusion. In other words, the ‘bottleneck’ in the mass transfer is due to the turbulent mixing and not due to the (rapid) cooling. This implies that resolving the mixing at the largest scales ($\sim R_{\rm cl}$) is sufficient to converge in the mass growth rate $\dot{m}_{\rm cl}$. This has been explicitly checked with the above mentioned high-resolution simulation which partially also included thermal conduction (Tan et al., 2021). The small increase in cold mass $m_{\rm cl}$ with increasing resolution is unsurprising as well, because we are able to capture the formation and evolution of smaller cold clumps when the resolution is higher. Thus, the overall cold gas surface area formed by radiative cooling, can be higher in these simulations – which is a major difference to the turbulent mixing layer simulations where the (effective) surface area is fixed. To which degree this additional fragmentation leads to a (non-)convergence in mass transfer depends on the exact geometry and other simulation parameters. Although it is an interesting avenue for future investigations, it does not affect our main conclusions here.
Interestingly, this slightly larger $\dot{m}_{\rm cl}$ for the simulations with $R_{\rm cl}/d_{\rm cell}|_{\rm ini}=8-32$ also leads to a higher transfer of momentum as seen in the position of the centre of mass of the cloud in the right panel of Fig. 23. Therefore, the cloud travels faster and gets quickly entrained at higher resolution. A similar trend has also been previously found in our earlier works on cloud-crushing (cf. Fig. B1 in Kanjilal et al. 2021). However, CC85 wind introduces an additional complication due to the faster travel of the cloud material – the wind condition depends strongly on the instantaneous position of the cloud (for example, $T_{\rm wind}\sim d^{-2(\gamma-1)}$), which might lead to differences in cloud properties across simulations at different resolutions as time progresses.
Fig. 24 shows the temporal evolution of the cloud ($T<2T_{\rm cl}=8\times 10^{4}\ \rm K$) profile from our fiducial simulation set (cf. Table 1). It is similar to Fig. 7, but at a resolution of $R_{\rm cl}/d_{\rm cell}|_{\rm ini}=16$; one level higher than the base fiducial resolution. The qualitative picture that clouds follow the local wind pressure as the they move downstream, and that they are slightly under-pressurized causing them to keep on growing (albeit gradually losing their density contrast, a.k.a ‘fading’) robustly carries over across simulations at different resolution. This demonstrates that our scaling arguments and the general description of cloud properties are not sensitive to the resolution chosen for the simulation.